# SEA: Self-Evolution Agent with Step-wise Reward for Computer Use

### v1

@@ III-F Temporal Compressed Sensing Mechanism
$\mathcal{I}_{\text{TCSM}}=\left\{\operatorname{Pad}\left(\operatorname{Resize}(I_{i})\right)\ \middle|\ i\in\mathcal{H},\ i\geq n-k+1\right\}$ (6)
@@ IV-A Experimental Setup
Implementation Details. For the training part, we use the LLaMa-Factory to train our grounding model which is used in the model integration. And for the whole model, we use the Easy-R1 to train the model by using the GRPO. During the whole training phase, we use one server with 8xA100(80GB) GPU. During the inference phase, we reference the GTA-1 [41] to have better results. Firstly, we used our agent to get $N$ possible results, in our experiment, $N$ equals to 8. And then we will use another model to select the best one from them, which is different from GTA-1 [41] using the GPT-o3.
@@
Datasets and Metrics. As we have mentioned before, we will use the generated verifiable tasks to produce the trajectory training data. In our experiments, we use the 400 tasks. And for each inference in the trajectory, we produce eight results. For the evaluation, we evaluate our improved grounding model on the ScreenSpot-V2 [23, 45] and ScreenSpotPro [53] datasets using accuracy as the evaluation metric. Subsequently, we evaluate our proposed model on the OSWorld benchmark [54], which is designed to evaluate the computer use agent’s performance on open-ended GUI tasks. This benchmark comprises 369 tasks spanning real-world web and desktop applications, providing a diverse and challenging platform to evaluate the agent’s ability to complete user tasks within a Linux environment. Performance is measured primarily by task success rate.
IV-B Experimental Results
IV-B1 Grounding Performance
@@
We first evaluate the performance of our improved and trained grounding model. Among models of comparable scale, our grounding model achieves SOTA performance, as shown in  I.
IV-B2 OS-World Performance
@@
We compare our method with state-of-the-art approaches on the OSWorld benchmark, with results summarized in II. As shown, despite utilizing significantly fewer parameters and incurring lower training costs, the proposed method still surpasses the larger model UI-TARS-72B-DPO [38] by 3.0% in task success rate. Furthermore, we reproduced the baseline model UI-TARS-1.5-7B [55] under the same experimental conditions. Compared to this baseline, SEA achieved a performance improvement of 5.2% and attained the highest task success rate among models of the same scale on the OSWorld benchmark, outperforming all existing state-of-the-art approaches. These results demonstrate the effectiveness of the proposed method in handling complex and realistic user tasks across diverse scenarios, highlighting the robustness and generalization capability of our approach. Meanwhile, as shown in this table, the model without the enhancement also achieves the SOTA performance, which proves the effectiveness of the data generation of verifiable tasks and the step-wise reinforcement learning.
TABLE II: Performance Comparison of the Proposed Method and other SOTA Models on the OSWorld Benchmark.
@@
OpenCUA-3B [56] 18.8
OpenCUA-Qwen2-7B [56] 23.1
OpenCUA-7B [56] 26.6
UI-TARS-72B-DPO [38] 27.1
UI-TARS-1.5-7B [55] 27.4 (24.9)
@@
IV-C Ablation Study
Generalization Enhancement on planning. As we have mentioned, we propose the enhancement method to improve the generalization of the model trained by RL. As presented in  II, the task completion rate increases from 28.1% to 30.1%, representing an improvement of 2.0%. This indicates that by integrating weights from different task modules and adopting a self-evolution mechanism, the model can achieve better generalization to novel task scenarios. We think that this kind of enhancement makes the original model have more knowledge of the PC’s screenshots, which will help the model produce more precise planning of different software, even if it is a new application.
V Conclusion
In this paper, we propose a series of innovations to produce a self-evolution for automatically operating the computers following the user’s instructions. We have proposed a novel pipeline to generate closed-loop verifiable tasks’ trajectories for training agents’ ability in reasoning, which solves the challenge of letting the model evolve itself from the data. In this part, we use two agents to generate the task instruction and the verification program, respectively. And then the generated verifiable tasks will be passed to extract the trajectory with assessment, which periodically replaces the generation model with the latest trained model and produces high-quality data. After this, the filter model will select the data for training. Secondly, we have proposed a step-wise reinforcement learning method to replace the long-horizon training, which frees the model to evolve from the computation limitation. Thirdly, we have proposed a enhancement method to help the model have the great grounding and planning ability at the time.
Experiments on the OSWorld benchmark demonstrate that the proposed agent achieves competitive results, which outperform other agents of the same scale and have comparable results to larger models. And we hope that our proposed method, including the pipeline to generate verifiable task trajectories, trajectory reasoning by step- wise reinforcement learning and grounding-based generalization enhancement can offer a promising new direction for the computer use agent area.

### v2

@@ 3.3.3. Temporal Compressed Sensing Mechanism
(6) $\mathcal{I}_{\text{TCSM}}=\left\{\operatorname{Pad}\left(\operatorname{Resize}(I_{i})\right)\ \middle|\ i\in\mathcal{H},\ i\geq n-k+1\right\}$
@@ 4.1. Experimental Setup
Implementation Details. For the training part, we use the LLaMa-Factory to train our grounding model, which is used in the model integration. And for the whole model, we use the Easy-R1 to train the model by using the GRPO. During the entire training phase, we use one server with 8 A100 (80GB) GPUs. During the inference phase, we reference the GTA-1 (Yang et al., 2025) to have better results. Firstly, we used our agent to get $N$ possible results; in our experiment, $N$ equals 8. And then we will use another model to select the best one from them, which differs from GTA-1 (Yang et al., 2025) in that it utilizes GPT-o3.
@@
Datasets and Metrics. As we have mentioned before, we will use the generated verifiable tasks to produce the trajectory training data. In our experiments, we use the 400 tasks. And for each inference in the trajectory, we produce eight results. For the evaluation, we evaluate our improved grounding model on the ScreenSpot-V2 (Kanzhi et al., 2024; Zhiyong et al., 2024b) and ScreenSpotPro (Li et al., 2025a) datasets using accuracy as the evaluation metric. Subsequently, we evaluate our proposed model on the OSWorld benchmark (Tianbao et al., 2024), which is designed to evaluate the computer use agent’s performance on open-ended GUI tasks. This benchmark comprises 369 tasks spanning real-world web and desktop applications, providing a diverse and challenging platform to evaluate the agent’s ability to complete user tasks within a Linux environment. Performance is measured primarily by task success rate.
4.2. Inference Stage
During the inference stage, we build upon the GTA1 (Yang et al., 2025) and replace the original inference model with our proposed model SEA. As a result, both the sampling and selection of candidate trajectories during inference are handled entirely by SEA, without relying on any external scoring models. This design significantly reduces computational overhead and effectively improves decision-making efficiency.
4.3. Experimental Results
4.3.1. Grounding Performance
@@
We first evaluate the performance of our improved and trained grounding model. Among models of comparable scale, our grounding model achieves SOTA performance, as shown in  Table 1. As shown in  Table 1, our grounding model has each domain in the ScreenSpotPro (Li et al., 2025a) and outperforms the larger models such as CogAgent-18B (Hong et al., 2024), Claude Computer Use (Hu et al., ) and UI-TARS-72B (Yujia et al., 2025). As ScreenSpotPro (Li et al., 2025a) does not explicitly reflect our model’s performance in the web domain, we also evaluate our model in the ScreenSpot-V2 (Kanzhi et al., 2024; Zhiyong et al., 2024b), shown in the  Table 2.
4.3.2. OS-World Performance
@@
We compare our method with state-of-the-art approaches on the OSWorld benchmark, with results summarized in Table 3. As shown, despite utilizing significantly fewer parameters and incurring lower training costs, the proposed method still surpasses the larger model UI-TARS-72B-DPO (Yujia et al., 2025) by 3.0% in task success rate. Furthermore, we reproduced the baseline model UI-TARS-1.5-7B (Seed, 2025) under the same experimental conditions. Compared to this baseline, SEA achieved a performance improvement of 5.2% and attained the highest task success rate among models of the same scale on the OSWorld benchmark, outperforming all existing state-of-the-art approaches. These results demonstrate the effectiveness of the proposed method in handling complex and realistic user tasks across diverse scenarios, highlighting the robustness and generalization capability of our approach. Meanwhile, as shown in this table, the model without the enhancement also achieves the SOTA performance, which proves the effectiveness of the data generation of verifiable tasks and the step-wise reinforcement learning.
Table 4. The offline results for the ablation study. In order to conveniently compare, we have evaluated some SOTA methods on our deployed OSWorld environment. We choose some representative methods, and the UI-TARS-1.5-7B is our foundation model.
@@
OpenCUA-3B (Wang et al., 2025) 18.8
OpenCUA-Qwen2-7B (Wang et al., 2025) 23.1
OpenCUA-7B (Wang et al., 2025) 26.6
UI-TARS-72B-DPO (Yujia et al., 2025) 27.1
UI-TARS-1.5-7B (Seed, 2025) 24.9
Ours(w/o merging + w/o enhancement) 27.6
@@
5. Ablation Study
In this section, we will figure out the function of modules in the proposed method through experiments. Firstly, we will demonstrate the effect of merging the grounding and planning models, and secondly, we will show whether the Temporal Compressed Sensing Mechanism enhances the model’s performance.
5.1. Merging models
As we have mentioned before, we have trained another grounding model to enhance the model’s ability to recognize screenshot information. Different from the previous methods using the two-stage training or distill models, we use the model merging method to make the final model have both the ability to extract information from the desktop and plan. As shown in  Table 4, before merging, our model achieves 27.6 in our offline evaluation. The reason why we use the offline instead of submitting the ablation experiment to the official online test is that we need to do many evaluations during the whole process. Meanwhile, we have submitted the final model to the official website and the result is shown in  Table 3. In order to have a comparable baseline, we have test the other SOTA methods in our offline evaluation. Comparing the baseline, the merging can improve our model’s accuracy from $27.6~\%$ to $28.1~\%$ for operating computers. We believe that the merging of grounding can enhance the planning model’s ability to understand screen images. During the process of using the proposed reinforcement learning to train models, the distribution of weights shifts to improve their ability to reason. So, when we merge the grounding model’s weights from layer to layer, it can strike a balance between reasoning and understanding the desktop images, and in other words, achieve better performance in the overall metric. This indicates that by integrating weights from different task modules and adopting a self-evolution mechanism, the model can achieve better generalization to novel task scenarios. We believe that this type of enhancement enables the original model to gain a deeper understanding of the PC’s screenshots, which will help the model produce more precise plans for various software, including new applications.
5.2. Generalization Enhancement on planning.
Besides the merging process, we also proposed using the Temporal Compressed Sensing Mechanism(TCSM) to enhance the final model. During our experiments, we found that TCSM is a simple but effective method to improve the model’s performance. As shown in  Table 4, the TCSM can improve the model from 28.1 to 30.1 without extra training and inference cost. We think the reason why it can improve the performance is that the temporal compression of the image is a kind of compression of the historical observation space, which can make the model focus on the current information and invoke the model’s reasoning more. Although the history, including the previous actions and the corresponding results, can make the model have the knowledge of what has happened, the superfluous information will hinder the model from making a correct judgment. Nevertheless, the TCSM can strike a balance between using historical knowledge and reasoning.
6. Conclusion
In this paper, we propose a suite of innovations to develop the Self-Evolution Agent (SEA) for autonomous computer operation, addressing key challenges in data scalability, long-horizon training, and multi-capability integration.
First, we design a closed-loop pipeline to generate verifiable task trajectories: a Task Agent generates executable instructions, while a Coding Agent synthesizes execution/verification programs. The GATE method further refines trajectories by prioritizing short, successful paths and filtering redundant steps. Second, we introduce Trajectory Reasoning by Step-wise Reinforcement Learning (TR-SRL), which replaces monolithic long-horizon training with step-wise rewards (step success, thought-action consistency, action format), enabling efficient learning. Third, we propose a Grounding-Based Generalization Enhancement method: merging grounding and planning models via DARE and optimizing perception with the Temporal Compressed Sensing Mechanism (TCSM), achieving dual capabilities without extra training.
Experimental results on the OSWorld benchmark show that SEA (7B parameters) outperforms all same-scale models (e.g., 5.2% improvement over UI-TARS-1.5-7B) and matches the performance of larger models (e.g., UI-TARS-72B-DPO). Future work will focus on extending SEA to complex multi-application scenarios and further reducing training/inference latency for edge-device deployment.